RSS Feed

Node n9-b15-lt1 malfunction

Start on: 2017-05-12 00:10:00 Status: Resolved Fault / Issue Finish on: 2017-05-12 02:41:00

Please be advised that at this moment we have issues with node n9-b15-lt1. Due to this temporary issues on servers in this nodes can occur.

UPDATE: Node was replaced with new one. Servers has started, everything should run just fine.

We apologize for temporary inconvenience.

Related servers / services / nodes

* n9-b15-lt1

Node n10-b11-lt1 maintenance

Start on: 2017-05-09 22:00:00 Status: Completed Maintenance Finish on: 2017-05-09 23:00:00

Please take a note, that on 2017-05-09 from 22:00 till 23:00 (GMT +3) we will perform maintenance during which cooler will be changed. This hardware replacement is necessary, in order to avoid node failure in the future.

Due to it all servers in this node will be rebooted, expect downtime no more than 15 minutes.

Related servers / services / nodes

* n10-b11-lt1

Node n16-b12-lt1 maintenance

Start on: 2017-05-02 23:00:00 Status: Completed Maintenance Finish on: 2017-05-03 01:00:00

Please take a note, that on 2017-05-02 from 23:00 till 2017-05-03 01:00 (GMT +2) we will perform maintenance during which RAM module will be changed. This hardware replacement is necessary, in order to avoid node failure due to malfunctioned RAM module.

Due to it all servers in this node will be rebooted, expect downtime no more than 15 minutes.

Node affected:

* n16-b12-lt1

Related servers / services / nodes